# UnMix-NeRF: Spectral Unmixing Meets Neural Radiance Fields

### v1

@@ 1 Introduction
Neural Radiance Fields (NeRF) [27] have revolutionized 3D scene representation, enabling photorealistic novel view synthesis through implicit volumetric rendering. Recent advances have extended NeRF to other tasks such as segmentation[6], object detection[12], and spatial reasoning [13], leveraging pre-trained vision models such as SAM [16] and CLIP [32] for higher-level scene understanding. Among these, semantic segmentation has emerged as a powerful tool, enabling the grouping of regions into object-level categories in multi-view settings [26]. A more challenging task lies in understanding and segmenting materials, moving beyond recognizing what an object is, to understanding what it is composed of. Materials are fundamental to how we interpret and interact with the world, as they define surface properties, functionality, and usability, enabling richer scene understanding and making them essential for applications such as robot autonomy [1], augmented reality [44], and simulation [37], where accurate material perception is crucial for decision-making and interaction.
@@
Spectral imaging (SI) emerges as a powerful framework for scene representation, as it provides rich spectral information across multiple wavelengths. This, in turn, allows for a more complete characterization of a scene, effectively encoding unique spectral cues for different materials. Integrating spectral data into implicit neural representations remains a challenge. Standard NeRF formulations model scene appearance using only RGB values, which inherently limits their capacity to differentiate materials with similar color but distinct spectral properties. While recent efforts have attempted to extend NeRF to spectral data [22, 3] by replacing the standard RGB output with per-channel spectral radiance predictions, these approaches treat spectral information as an additional dimension in the output space without explicitly leveraging the inherent structure of it, e.g. materials have representative spectral signatures.
@@
Contributions. In this paper we propose UnMix-NeRF, the first framework that integrates spectral unmixing into NeRF, allowing joint spectral novel view synthesis and unsupervised material segmentation. Our method learns global endmembers through a dictionary optimized during training and models spectral reflectance via a diffuse and specular decomposition. The diffuse component is represented by per-point abundances and scaling factors, while a dedicated branch predicts view-dependent specular effects. The final spectral radiance is obtained by combining both components. Moreover, the per-point spectral unmixing naturally enables unsupervised material segmentation by leveraging the rendered abundance vectors. To summarize, our key contributions are:
@@
Our approach not only outperforms state-of-the-art methods in hyperspectral view synthesis, but also achieves pixel-accurate 3D material segmentation. Also, we extend a synthetic hyperspectral dataset with paired 3D material annotations, enabling rigorous evaluation of spectral unmixing and segmentation in complex scenes.
@@ 2 Related Work
Neural Radiance Fields. NeRF revolutionized novel view synthesis by representing scenes as continuous neural fields. Through volumetric rendering and multi-view optimization, it learns to predict density and color at any 3D point and viewing direction, enabling photorealistic view synthesis [27]. Several works have extended NeRF to new imaging modalities: SeaThru-NeRF [19] models underwater light scattering, Thermal-NeRF [41] handles infrared emissions, LiDAR-NeRF [35] enables LiDAR point cloud synthesis, and NeSpoF [15] incorporates polarization information. For hyperspectral imaging, HS-NeRF [3] and Spec-NeRF [20] directly extend NeRF to predict per-wavelength radiance. Recent approaches like SpectralNeRF [22], introduces physically-based spectral rendering by extending NeRF with an additional autoencoder to predict per-wavelength radiance, and HyperGS [36], a concurrent work that adapts 3D Gaussian Splatting for spectral scene reconstruction. However, these methods treat spectral bands as independent channels, failing to take advantage of the rich material information encoded within the spectral signatures. These extensions require substantial computational overhead and also ignore the structured nature of spectral information, limiting their ability to disentangle materials.
Spectral Unmixing. Spectral unmixing decomposes pixels into endmembers (materials signatures) and their abundances (percentage of material per pixel) [14]. Beyond its use in remote sensing, unmixing enables material understanding in indoor and outdoor scenes in robotics [5], industrial inspection [8], and material classification [43]. The process typically involves identifying the different concurrent endmember spectra and their fractional abundances, per pixel, while enforcing physical constraints like non-negativity and sum-to-one (i.e. the sum of all abundances must be 100%). Classical approaches employ the linear mixing model (LMM), which assumes linear combination of endmembers [14]. However, LMM struggles with spectral variability caused by illumination, atmospheric conditions, and material mixing effects [2]. Extensions like the Augmented LMM (ALMM) [11] and the Extended LMM (ELMM)[38] address this by incorporating scaling factors and environmental variations through learned dictionaries and bundle-based representations, improving abundance estimation accuracy under real-world conditions. Recent deep learning approaches leverage transformers and convolutional neural networks (CNNs) to learn endmembers and abundances directly from data [9, 42]. While these methods have shown promise for 2D spectral unmixing, they are currently limited to single-view analysis and have not been extended to 3D reconstruction, where material decomposition could enable richer scene understanding. The challenge of incorporating physical unmixing constraints into neural architectures for 3D scenes remains unexplored.
@@ 3.2 Spectral Unmixing
ensuring that abundances are non-negative, and that they sum to one, enforcing a physically meaningful representation of spectral mixing.
@@ 4.1 Spectral Unmixing Field
thus ensuring non-negativity and the sum-to-one constraint according to Eq. (4). Additionally, $h$ is passed through a sigmoid activation to constraint its values within $[0,1]$, thus modulating the contribution of the diffuse color. Each abundance $\tilde{a}_{k}$ corresponds to the fractional presence of the $k$-th material at point $\mathbf{x}$, providing a physically meaningful decomposition of the current spectral signature. With these, the diffuse reflectance is computed as
@@ 4.5 Loss Function
with hyperparameters $\lambda_{\mathrm{spec}}$ and $\lambda_{\mathrm{rgb}}$ balancing the importance of hyperspectral versus RGB fidelity. Minimizing this loss does not only leads to accurate novel view synthesis, but also ensures physically consistent spectral unmixing. This results in both high-quality spectral reconstructions and reliable material segmentation through the per-ray abundance predictions.
@@ 4.6 Implementation Details
We implemented our method within the Nerfstudio framework [34], building on top of the Nerfacto implementation. The hidden feature dimension $D$ was set to 16, and we integrated Nerfacc [21] into the volumetric rendering and sampling process to reduce training time. Additionally, we applied gradient scaling based on the squared ray distance to each pixel, as suggested in [31]. For optimization, we used the Adam optimizer with a learning rate of $1\times 10^{-2}$ and an exponential learning rate scheduler. All models were trained for $20,000$ iterations and all experiments were performed on an NVIDIA A100 GPU, with a maximum memory usage of approximately 20 GB. For all experiments, we set $\lambda_{\mathrm{spec}}=5$ and $\lambda_{\mathrm{rgb}}=1$. To ensure physically meaningful spectral signatures, the endmembers $\mathbf{E}$ was constrained at each iteration by applying element-wise clamping within the range $[0,1]$. For each trained scene, the number of endmembers $K$ is selected as a heuristic approximation of the number of distinct materials present in the scene.
@@ 5.1 Datasets, Preprocessing, Metrics and Baselines
Metrics. We evaluate UnMix-NeRF and competing methods on hyperspectral view synthesis using rigorous metrics that capture both accuracy and spectral fidelity, including PSNR, SSIM, Spectral Angle Mapping (SAM), and RMSE. For unsupervised material segmentation, we report standard metrics, including mean Intersection over Union (mIoU) and F1 score.
@@ 5.2 Ablation Studies
Results in Table 1 illustrate the incremental contributions of each component added to our UnMix-NeRF baseline. The initial model (Unmix Only) achieves moderate spectral reconstruction quality (PSNR $=24.73$), establishing a performance baseline without additional constraints. Introducing physical constraints reduces the PSNR slightly by a margin of $-2.37$ due to the restrictive nature of these constraints on spectral variability. The addition of scaling factors significantly boosts the performance (PSNR increases by $+7.00$, SAM decreases by $-0.049$, and RMSE decreases by $-0.042$), underscoring their importance in modeling local spectral variations. Incorporating the RGB loss further improves the spectral accuracy (PSNR improves by $+1.56$), demonstrating the benefit of coupling hyperspectral predictions with RGB supervision. Using the VCA initialization provides a minor improvement of $+0.17$ in PSNR, indicating that an informed initialization of the endmembers aids the optimization stability. Finally, integrating the specular field yields the best overall results (PSNR increases by $+2.11$, SSIM reaches $0.935$), confirming that explicitly modeling specular reflectance substantially enhances spectral rendering and overall image quality.
@@ 5.4 Quantitative Results
Table 3: Quantitative results on the BaySpec dataset for the Pinecone, Caladium, and Anacampseros scenes. Our method achieves the best performance across all metrics
@@
We further evaluate UnMix-NeRF on the BaySpec dataset, as shown in Table 3. Our method consistently outperforms all baselines in all scenes, achieving the highest PSNR, SSIM, and the lowest SAM and RMSE values. Notably, our model improves over HyperGS by $+3.15$ PSNR on the Caladium scene. Similar trends are observed in the Pinecone and Anacampseros scenes, where our approach provides both higher reconstruction fidelity and better spectral consistency. These results show that explicitly modeling spectral unmixing in NeRF improves spectral prediction accuracy over per-pixel regression.
Table 4 reports the results on the Surface Optics dataset. Our method achieves the best performance in both the Rosemary and Basil scenes, obtaining a $+2.14$ and $+3.91$ PSNR improvement over HyperGS, respectively. Furthermore, our approach achieves the lowest SAM and RMSE values, confirming its ability to accurately reconstruct spectral details across diverse material compositions. The results for the remaining scenes are provided in the supplementary material.
@@ 6 Conclusions
Acknowledgments. This work is supported by the KAUST Center of Excellence for Generative AI under award number 5940. The computational resources are provided by IBEX, which is managed by the Supercomputing Core Laboratory at KAUST.

### v2

@@ 1 Introduction
Neural Radiance Fields (NeRF) [31] have revolutionized 3D scene representation, enabling photorealistic novel view synthesis through implicit volumetric rendering. Recent advances have extended NeRF to other tasks such as segmentation [7], object detection [14], and spatial reasoning [16], leveraging pre-trained vision models such as SAM [19] and CLIP [36] for higher-level scene understanding. Among these, semantic segmentation has emerged as a powerful tool, enabling the grouping of regions into object-level categories in multi-view settings [30]. A more challenging task lies in understanding and segmenting materials, moving beyond recognizing what an object is to understanding what it is composed of. Materials are fundamental to how we interpret and interact with the world, as they define surface properties, functionality, and usability, enabling richer scene understanding and making them essential for applications such as robot autonomy [1], augmented reality [48], and simulation [41], where accurate material perception is crucial for decision-making and interaction.
@@
Spectral imaging (SI) emerges as a powerful framework for scene representation, as it provides rich spectral information across multiple wavelengths [15, 12, 5]. This, in turn, allows for a more complete characterization of a scene, effectively encoding unique spectral cues for different materials. Integrating spectral data into implicit neural representations remains a challenge. Standard NeRF formulations model scene appearance using only RGB values, which inherently limits their capacity to differentiate materials with similar color but distinct spectral properties. While recent efforts have attempted to extend NeRF to spectral data [26, 3] by replacing the standard RGB output with per-channel spectral radiance predictions, these approaches treat spectral information as an additional dimension in the output space without explicitly leveraging its inherent structure, e.g. materials have representative spectral signatures.
@@
Contributions. In this paper, we propose UnMix-NeRF, the first framework that integrates spectral unmixing into NeRF, allowing joint spectral novel view synthesis and unsupervised material segmentation. Our method learns global endmembers through a dictionary optimized during training and models spectral reflectance via a diffuse and specular decomposition. The diffuse component is represented by per-point abundances and scaling factors, while a dedicated branch predicts view-dependent specular effects. The final spectral radiance is obtained by combining both components. Moreover, the per-point spectral unmixing naturally enables unsupervised material segmentation by leveraging the rendered abundance vectors. To summarize, our key contributions are:
  • (i)
@@
Our approach not only outperforms state-of-the-art methods in hyperspectral view synthesis but also achieves pixel-accurate 3D material segmentation. Also, we extend a synthetic hyperspectral dataset with paired 3D material annotations, enabling rigorous evaluation of spectral unmixing and segmentation in complex scenes.
@@ 2 Related Work
Neural Radiance Fields. NeRF revolutionized novel view synthesis by representing scenes as continuous neural fields. Through volumetric rendering and multi-view optimization, it learns to predict density and color at any 3D point and viewing direction, enabling photorealistic view synthesis [31]. Several works have extended NeRF to new imaging modalities: SeaThru-NeRF [23] models underwater light scattering, Thermal-NeRF [45] handles infrared emissions, LiDAR-NeRF [39] enables LiDAR point cloud synthesis, and NeSpoF [18] incorporates polarization information. For hyperspectral imaging, HS-NeRF [3] and Spec-NeRF [24] directly extend NeRF to predict per-wavelength radiance. Recent approaches like SpectralNeRF [26], introduce physically-based spectral rendering by extending NeRF with an additional autoencoder to predict per-wavelength radiance, and HyperGS [40], a concurrent work that adapts 3D Gaussian Splatting for spectral scene reconstruction. However, these methods treat spectral bands as independent channels, failing to take advantage of the rich material information encoded within the spectral signatures. These extensions require substantial computational overhead and also ignore the structured nature of spectral information, limiting their ability to disentangle materials.
Spectral Unmixing. Spectral unmixing decomposes pixels into endmembers (materials signatures) and their abundances (percentage of material per pixel) [17]. Beyond its use in remote sensing, unmixing enables material understanding in indoor and outdoor scenes in robotics [6], industrial inspection [9], and material classification [47]. The process typically involves identifying the different concurrent endmember spectra and their fractional abundances, per pixel, while enforcing physical constraints like non-negativity and sum-to-one (i.e., the sum of all abundances must be 100%). Classical approaches employ the linear mixing model (LMM), which assumes a linear combination of endmembers [17]. However, LMM struggles with spectral variability caused by illumination, atmospheric conditions, and material mixing effects [2]. Extensions like the Augmented LMM (ALMM) [13] and the Extended LMM (ELMM) [42] address this by incorporating scaling factors and environmental variations through learned dictionaries and bundle-based representations, improving abundance estimation accuracy under real-world conditions. Recent deep learning approaches leverage transformers and convolutional neural networks (CNNs) to learn endmembers and abundances directly from data [10, 46]. While these methods have shown promise for 2D spectral unmixing, they are currently limited to single-view analysis and have not been extended to 3D reconstruction, where material decomposition could enable richer scene understanding. The challenge of incorporating physical unmixing constraints into neural architectures for 3D scenes remains unexplored.
@@ 3.2 Spectral Unmixing
ensuring that abundances are non-negative and that they sum to one, enforcing a physically meaningful representation of spectral mixing.
@@ 4.1 Spectral Unmixing Field
thus ensuring non-negativity and the sum-to-one constraint according to Eq. (4). Additionally, $h$ is passed through a sigmoid activation to constrain its values within $[0,1]$, thus modulating the contribution of the diffuse color. Each abundance $\tilde{a}_{k}$ corresponds to the fractional presence of the $k$-th material at point $\mathbf{x}$, providing a physically meaningful decomposition of the current spectral signature. With these, the diffuse reflectance is computed as follows:
@@ 4.5 Loss Function
with hyperparameters $\lambda_{\mathrm{spec}}$ and $\lambda_{\mathrm{rgb}}$ balancing the importance of hyperspectral versus RGB fidelity. Minimizing this loss not only leads to accurate novel view synthesis but also ensures physically consistent spectral unmixing. This results in both high-quality spectral reconstructions and reliable material segmentation through the per-ray abundance predictions.
@@ 4.6 Implementation Details
We implemented our method within the Nerfstudio framework [38], building on top of the Nerfacto implementation. The hidden feature dimension $D$ was set to 16, and we integrated Nerfacc [25] into the volumetric rendering and sampling process to reduce training time. Additionally, we applied gradient scaling based on the squared ray distance to each pixel, as suggested in [35]. For optimization, we used the Adam optimizer with a learning rate of $1\times 10^{-2}$ and an exponential learning rate scheduler. All models were trained for $20,000$ iterations, and all experiments were performed on an NVIDIA A100 GPU, with a maximum memory usage of approximately 20 GB. For all experiments, we set $\lambda_{\mathrm{spec}}=5$ and $\lambda_{\mathrm{rgb}}=1$. To ensure physically meaningful spectral signatures, the endmembers $\mathbf{E}$ were constrained at each iteration by applying element-wise clamping within the range $[0,1]$. For each trained scene, the number of endmembers $K$ is selected as a heuristic approximation of the number of distinct materials present in the scene.
@@ 5.1 Datasets, Preprocessing, Metrics and Baselines
Metrics. We evaluate UnMix-NeRF and competing methods on hyperspectral view synthesis using rigorous metrics that capture both accuracy and spectral fidelity, including PSNR, SSIM, Spectral Angle Mapping (SAM) [20], and RMSE. For unsupervised material segmentation, we report standard metrics, including mean Intersection over Union (mIoU) and F1 score.
@@ 5.2 Ablation Studies
Results in Table 1 illustrate the incremental contributions of each component added to our UnMix-NeRF baseline. The initial model (Unmix Only) achieves moderate spectral reconstruction quality (PSNR =24.73=24.73), establishing a performance baseline without additional constraints. Introducing physical constraints reduces the PSNR slightly by a margin of 2.37-2.37 due to the restrictive nature of these constraints on spectral variability. The addition of scaling factors significantly boosts the performance (PSNR increases by +7.00+7.00, SAM decreases by 0.049-0.049, and RMSE decreases by 0.042-0.042), underscoring their importance in modeling local spectral variations. Incorporating the RGB loss further improves the spectral accuracy (PSNR improves by +1.56+1.56), demonstrating the benefit of coupling hyperspectral predictions with RGB supervision. Using the VCA initialization provides a minor improvement of +0.17+0.17 in PSNR, indicating that an informed initialization of the endmembers aids the optimization stability. Finally, integrating the specular field yields the best overall results (PSNR increases by +2.11+2.11, SSIM reaches $0.935$), confirming that explicitly modeling specular reflectance substantially enhances spectral rendering and overall image quality.
@@ 5.4 Quantitative Results
Table 3: Quantitative results on the BaySpec dataset for the Pinecone, Caladium, and Anacampseros scenes. Our method achieves the best performance across all metrics.
@@
We further evaluate UnMix-NeRF on the BaySpec dataset, as shown in Table 3. Our method consistently outperforms all baselines in all scenes, achieving the highest PSNR, SSIM, and the lowest SAM and RMSE values. Notably, our model improves over HyperGS by +3.15+3.15 PSNR on the Caladium scene. Similar trends are observed in the Pinecone and Anacampseros scenes, where our approach provides both higher reconstruction fidelity and better spectral consistency. These results show that explicitly modeling spectral unmixing in NeRF improves spectral prediction accuracy over per-pixel regression.
Table 4 reports the results on the Surface Optics dataset. Our method achieves the best performance in both the Rosemary and Basil scenes, obtaining a +2.14+2.14 and +3.91+3.91 PSNR improvement over HyperGS, respectively. Furthermore, our approach achieves the lowest SAM and RMSE values, confirming its ability to accurately reconstruct spectral details across diverse material compositions. The results for the remaining scenes are provided in the supplementary material.
@@ 6 Conclusions
Acknowledgments. The research reported in this publication was supported by funding from King Abdullah University of Science and Technology (KAUST) - Center of Excellence for Generative AI, under award number 5940.